# BACON: A fully explainable AI model with graded logic for decision making problems

### v2

@@ G.2 Pruning the binary tree
| 1 | fractal_dimension3 | smoothness2 | 0.6225 | 1.7061 | 0.3775 | 98.42% |
| 2 | Node1 | texture2 | 0.6225 | 1.3767 | 0.3775 | 98.42% |
| 3 | Node2 | perimeter1 | 0.6225 | -0.1532 | 0.3775 | 98.42% |
| 4 | Node3 | compactness1 | 0.6225 | 1.5872 | 0.3775 | 98.42% |
| 5 | Node4 | smoothness1 | 0.6225 | 0.4842 | 0.3775 | 98.42% |
| 6 | Node5 | symmetry1 | 0.6225 | 1.8122 | 0.3775 | 98.42% |
| 7 | Node6 | fractal_dimension2 | 0.6225 | 1.5481 | 0.3775 | 98.42% |
| 8 | Node7 | texture1 | 0.6225 | 1.5347 | 0.3775 | 98.42% |
| 9 | Node8 | compactness3 | 0.6225 | 1.6550 | 0.3775 | 98.42% |
| 10 | Node9 | fractal_dimension1 | 0.6225 | 1.5337 | 0.3775 | 98.42% |
| 11 | Node10 | concavity2 | 0.6225 | -0.7101 | 0.3775 | 98.42% |
| 12 | Node11 | concave_points2 | 0.6225 | -0.6050 | 0.3775 | 98.42% |
| 13 | Node12 | compactness2 | 0.6225 | -0.8014 | 0.3775 | 98.42% |
| 14 | Node13 | symmetry2 | 0.6225 | 1.9969 | 0.3775 | 98.42% |
| 15 | Node14 | concavity1 | 0.6225 | 1.9263 | 0.3775 | 98.42% |
| 16 | Node15 | area3 | 0.6226 | -0.7999 | 0.3774 | 98.42% |
| 17 | Node16 | smoothness3 | 0.7365 | 1.6205 | 0.2635 | 91.04% |
| 18 | Node17 | concave_points1 | 0.7979 | -0.6530 | 0.2021 | 91.04% |
| 19 | Node18 | concave_points3 | 0.7909 | 1.8086 | 0.2091 | 87.70% |
| 20 | Node19 | area2 | 0.7819 | -0.9105 | 0.2181 | 87.70% |
| 21 | Node20 | radius2 | 0.6848 | 0.2440 | 0.3152 | 85.59% |
| 22 | Node21 | texture3 | 0.8521 | 1.9154 | 0.1479 | 85.59% |
| 23 | Node22 | perimeter3 | 0.8782 | -0.8904 | 0.1218 | 85.59% |
| 24 | Node23 | concavity3 | 0.9171 | 1.9449 | 0.0829 | 84.53% |
| 25 | Node24 | radius3 | 0.9096 | -0.9351 | 0.0904 | 84.53% |
| 26 | Node25 | area1 | 0.8596 | 0.2491 | 0.1404 | 63.09% |
| 27 | Node26 | symmetry3 | 0.8951 | 1.8458 | 0.1049 | 63.09% |
| 28 | Node27 | perimeter2 | 0.9730 | -0.9525 | 0.0270 | 63.09% |
| 29 | Node28 | radius1 | 0.9974 | 1.5575 | 0.0026 | 62.74% |
@@ G.4 Multiple explanations
| 1 | concave_points3 | radius2 | symmetry3 | smoothness3 | area2 | 97.36% |
| 2 | radius2 | area2 | perimeter3 | concavity2 | compactness1 | 97.72% |
| 3 | symmetry2 | area2 | smoothness3 | fractal_dimension3 | concavity1 | 97.89% |
| 4 | radius2 | perimeter3 | smoothness3 | symmetry3 | concavity1 | 97.89% |
| 5 | area3 | smoothness3 | perimeter3 | perimeter2 | perimeter1 | 98.07% |
| 6 | radius3 | area3 | smoothness3 | concavity1 | symmetry3 | 97.72% |
| 7 | area3 | concave_points3 | radius2 | texture1 | smoothness3 | 98.24% |
| 8 | smoothness3 | concavity3 | area3 | radius1 | area2 | 98.24% |
| 9 | radius3 | texture3 | concavity1 | concavity2 | symmetry3 | 98.42% |
| 10 | radius3 | area3 | concavity_points1 | area2 | perimeter3 | 98.95% |
@@ Appendix H LLM generated reports
System
You are a report generator that produces a precise, human-readable,
and insightful explanation of an LSP aggregation tree. Use your
knowledge of LSP as well as the provided context to generate a
clear and logical report.
Context
This aggregation tree is for breast cancer diagnosis using the
WDSD dataset.
Background
LSP aggregators are defined in the following markdown table:
| GCD | Type | Subtype | Code | Name | Andness | Verbalization |
...(omitted)...
Instructions
1. Organize the report into the following sections:
- Overview: Summarize the overall decision logic of the tree.
- Decision Logic Walkthrough: Explain step by step from root
to leaves, introducing the operator, its meaning (use the
verbalization from the table where applicable), and how the
features are combined.
- Clinical Interpretation: Relate the aggregation logic to
known breast cancer diagnosis patterns (e.g., large size,
irregular shape), and highlight any interesting interactions
or compensations observed.
2. Use plain, human-friendly language suitable for clinicians or
decision-makers.
3. When describing operators, use both their code (e.g., "HHC")
and verbalization (e.g., "High Hyper-Conjunction extremely
strict ’must have all’ behavior").
4. Highlight the most influential features and explain how they
affect the overall decision.
5. Avoid excessive technical terms unless necessary, and prefer
clear analogies and explanations.
Input
The aggregation tree is expressed in the following JSON.
‘‘‘json
"operator": "HHC",
"children": [
"feature": "concave_points3",
"weight": 0.5,
"operator": "HHD",
"children": [
"feature": "radius2",
"weight": 0.04,
"feature": "area1",
"weight": 0.11
"feature": "concavity1",
"weight": 0.34
"feature": "perimeter3",
"weight": 0.51

### v3

@@ G.2 Pruning the binary tree
1 | fractal_dimension3 | smoothness2 | 0.6225 | 1.7061 | 0.3775 | 98.42% |
2 | Node1 | texture2 | 0.6225 | 1.3767 | 0.3775 | 98.42% |
3 | Node2 | perimeter1 | 0.6225 | -0.1532 | 0.3775 | 98.42% |
4 | Node3 | compactness1 | 0.6225 | 1.5872 | 0.3775 | 98.42% |
5 | Node4 | smoothness1 | 0.6225 | 0.4842 | 0.3775 | 98.42% |
6 | Node5 | symmetry1 | 0.6225 | 1.8122 | 0.3775 | 98.42% |
7 | Node6 | fractal_dimension2 | 0.6225 | 1.5481 | 0.3775 | 98.42% |
8 | Node7 | texture1 | 0.6225 | 1.5347 | 0.3775 | 98.42% |
9 | Node8 | compactness3 | 0.6225 | 1.6550 | 0.3775 | 98.42% |
10 | Node9 | fractal_dimension1 | 0.6225 | 1.5337 | 0.3775 | 98.42% |
11 | Node10 | concavity2 | 0.6225 | -0.7101 | 0.3775 | 98.42% |
12 | Node11 | concave_points2 | 0.6225 | -0.6050 | 0.3775 | 98.42% |
13 | Node12 | compactness2 | 0.6225 | -0.8014 | 0.3775 | 98.42% |
14 | Node13 | symmetry2 | 0.6225 | 1.9969 | 0.3775 | 98.42% |
15 | Node14 | concavity1 | 0.6225 | 1.9263 | 0.3775 | 98.42% |
16 | Node15 | area3 | 0.6226 | -0.7999 | 0.3774 | 98.42% |
17 | Node16 | smoothness3 | 0.7365 | 1.6205 | 0.2635 | 91.04% |
18 | Node17 | concave_points1 | 0.7979 | -0.6530 | 0.2021 | 91.04% |
19 | Node18 | concave_points3 | 0.7909 | 1.8086 | 0.2091 | 87.70% |
20 | Node19 | area2 | 0.7819 | -0.9105 | 0.2181 | 87.70% |
21 | Node20 | radius2 | 0.6848 | 0.2440 | 0.3152 | 85.59% |
22 | Node21 | texture3 | 0.8521 | 1.9154 | 0.1479 | 85.59% |
23 | Node22 | perimeter3 | 0.8782 | -0.8904 | 0.1218 | 85.59% |
24 | Node23 | concavity3 | 0.9171 | 1.9449 | 0.0829 | 84.53% |
25 | Node24 | radius3 | 0.9096 | -0.9351 | 0.0904 | 84.53% |
26 | Node25 | area1 | 0.8596 | 0.2491 | 0.1404 | 63.09% |
27 | Node26 | symmetry3 | 0.8951 | 1.8458 | 0.1049 | 63.09% |
28 | Node27 | perimeter2 | 0.9730 | -0.9525 | 0.0270 | 63.09% |
29 | Node28 | radius1 | 0.9974 | 1.5575 | 0.0026 | 62.74% |
@@ G.4 Multiple explanations
1 | concave_points3 | radius2 | symmetry3 | smoothness3 | area2 | 97.36% |
2 | radius2 | area2 | perimeter3 | concavity2 | compactness1 | 97.72% |
3 | symmetry2 | area2 | smoothness3 | fractal_dimension3 | concavity1 | 97.89% |
4 | radius2 | perimeter3 | smoothness3 | symmetry3 | concavity1 | 97.89% |
5 | area3 | smoothness3 | perimeter3 | perimeter2 | perimeter1 | 98.07% |
6 | radius3 | area3 | smoothness3 | concavity1 | symmetry3 | 97.72% |
7 | area3 | concave_points3 | radius2 | texture1 | smoothness3 | 98.24% |
8 | smoothness3 | concavity3 | area3 | radius1 | area2 | 98.24% |
9 | radius3 | texture3 | concavity1 | concavity2 | symmetry3 | 98.42% |
10 | radius3 | area3 | concavity_points1 | area2 | perimeter3 | 98.95% |
@@ Appendix H LLM generated reports
System ======= You are a report generator that produces a precise, human-readable, and insightful explanation of an LSP aggregation tree. Use your knowledge of LSP as well as the provided context to generate a clear and logical report. Context ======= This aggregation tree is for breast cancer diagnosis using the WDSD dataset. Background ========== LSP aggregators are defined in the following markdown table: | GCD | Type | Subtype | Code | Name | Andness | Verbalization | |--------|--------|--------|--------|--------|--------|--------| ...(omitted)... Instructions ============ 1. Organize the report into the following sections: - Overview: Summarize the overall decision logic of the tree. - Decision Logic Walkthrough: Explain step by step from root to leaves, introducing the operator, its meaning (use the verbalization from the table where applicable), and how the features are combined. - Clinical Interpretation: Relate the aggregation logic to known breast cancer diagnosis patterns (e.g., large size, irregular shape), and highlight any interesting interactions or compensations observed. 2. Use plain, human-friendly language suitable for clinicians or decision-makers. 3. When describing operators, use both their code (e.g., "HHC") and verbalization (e.g., "High Hyper-Conjunction extremely strict ’must have all’ behavior"). 4. Highlight the most influential features and explain how they affect the overall decision. 5. Avoid excessive technical terms unless necessary, and prefer clear analogies and explanations. Input ====== The aggregation tree is expressed in the following JSON. ‘‘‘json { "operator": "HHC", "children": [ { "feature": "concave_points3", "weight": 0.5, }, { "operator": "HHD", "children": [ { "feature": "radius2", "weight": 0.04, }, { "feature": "area1", "weight": 0.11 }, { "feature": "concavity1", "weight": 0.34 }, { "feature": "perimeter3", "weight": 0.51 } ] } ] } ‘‘‘